suppose that r varies directly as p and inversely as q^2, and that r = 27 when p =3 and q = 2. Find r when p = 2 and q = 3.
Answer by stanbon(48558)

suppose that r varies directly as p and inversely as q^2, and that r = 27 when p =3 and q = 2. Find r when p = 2 and q = 3.
Since r varies directly as p and inversely as q^2,
r=kp/q^2
Need to find "k"
Since r=27 when p=3 and q=2,
27=k(3)/2^2
27=3k/4
108=3k
k=36
Equation:
r=(36)p/q^2
Now find r when p=2 and q=3.
r=(36)2/9
r=8
Cheers,
Stan H.